OpenStudy (eric_d):

If (x-a) is a factor of the expression ax^3-3x^2-5ax-9, find the possible values of a. Factorise the expression for each of the values of a.

OpenStudy (eric_d):

@DLS

OpenStudy (eric_d):

@CGGURUMANJUNATH

OpenStudy (eric_d):

@aum

OpenStudy (callisto):

Let f(x) = ax^3-3x^2-5ax-9 Since (x-a) is a factor of f(x), f(a) = 0. Solve this equation to find the value of a first.

OpenStudy (eric_d):

a^4-8a^2-9=0 like this

OpenStudy (callisto):

Yes.

OpenStudy (eric_d):

what's nxt

OpenStudy (callisto):

Solve this equation

OpenStudy (eric_d):

I don't know hw to factorise this equation..

OpenStudy (callisto):

Let a^2 = b, then it becomes b^2-8b-9=0 Can you factorize the left side of the equation?

OpenStudy (eric_d):